Complete the balance sheet and sales information using the following financial data:

Total assets turnover: 1.3×
Days sales outstanding: 36.5 daysa
Inventory turnover ratio: 4×
Fixed assets turnover: 2.5×
Current ratio: 2.0×
Gross profit margin on sales: (Sales - Cost of goods sold)/Sales = 35%
aCalculation is based on a 365-day year.

Do not round intermediate calculations. Round your answers to the nearest dollar.

Balance Sheet
Cash $    Current liabilities $   
Accounts receivable     Long-term debt 72,000
Inventories     Common stock    
Fixed assets     Retained earnings 108,000
Total assets $360,000 Total liabilities and equity $   
Sales $    Cost of goods sold $   
Answer #1
Total asset turnover = Sales / Total assets
1.3 = Sales / 360000
Sales = 1.3 * 360000 468000
Days sales outstanding = ( Accounts receivables / Sales ) * Days in a year
36.5 = ( Accounts receivables / 468000 ) * 365
Accounts receivables = 36.5 * 468000 / 365 46800
Gross profit margin on sales = ( Sales - Cost of goods sold ) / Sales
35% = ( 468000 - Cost of goods sold ) / 468000 163800
35% * 468000 = 468000 - Cost of goods sold
163800 = 468000 - Cost of goods sold
Cost of goods sold = 468000 - 163800 304200
Inventory turnover ratio = Cost of goods sold / Inventories
4 = 304200 / Inventories
Inventories = 304200 / 4 76050
Fixed asset turnover = Sales / Fixed assets
2.5 = 468000 / Fixed assets
Fixed assets = 468000 / 2.5 187200
Cash = Total assets - Accounts receivables - Inventories - Fixed assets = 360000 - 46800 - 76050 - 187200 49950
Total liabilities and equity = Total assets 360000
Current assets = Cash + Accounts receivables + Inventories = 49950 + 46800 + 76050 172800
Current ratio = Current assets / Current liabilities
2 = 172800 / Current liabilities
Current liabilities = 172800 / 2 86400
Common stock = Total liabilities and equity - Current liabilities - Long term debt - Retained earnings = 360000 - 86400 - 72000 - 108000 93600
Cash 49950 Current liabilities 86400
Accounts receivables 46800 Long-term debt 72000
Inventories   76050 Common stock 93600
Fixed assets 187200 Retained earnings 108000
Total assets 360000 Total liabilities and equity 360000
Sales 468000 Cost of goods sold 304200